Method Substitution
12. Explain why the following linear system is not easy to solve by substitution.
3x+4y=10
2x-5y=9
Thank you very much and i need it today.

Answer by checkley77(12844)

3x+4y=10
3x=10-4y
x=(10-4y)/3
2x-5y=9
2(10-4y)/3-5y=9
(20-8y)/3-5y=9
(20-8y)/3=9+5y
20-8y=3(9+5y)
20-8y=27+15y
-8y-15y=27-20
-23y=7
y=7/-23
y=-7/23 ans.
2x-5(-7/23)=9
2x+35/23=9
2x=9-35/23
2x=172/23
x=172/23*1/2
x=172/46
x=86/23 ans.
Substitution sometimes create complex fractions which makes the process very difficult.